Soybean meal is 16% protein, cornmeal is 8% protein. How many pounds of each should be mixed togeather in order to get a 320 lb mixture that is 14% protein?
Soybean=
Coenmeal=

Answer by ptaylor(1894)

Soybean=240 lb
Coenmeal=80 lb
Let x=amount of 16% protein needed
Then 320-x=amount of 8% protein needed
Now we know that the amount of pure protein in the 16% mixture (0.16x) plus the amount of pure protein in the 8% mixture (0.08(320-x)) has to equal the amount of pure protein in the final mixture (0.14*320). So our equation to solve is:
0.16x+ 0.08(320-x)=0.14*320 get rid of parens and simplify
0.16x+25.6-0.08x=44.8 subtract 25.6 from each side
0.16x+25.6-25.6-0.08x=44.8-25.6 collect like terms
0.08x=19.2
divide each side by 0.08
x=240 lb-----------------------amount of 16% mixture needed(soybean)
320-x=320-240=80 lb-------------amount of 8% mixture needed(cornmeal)
CK
240*0.16+80*0.08=320*0.14
38.4+6.4=44.8
44.8=44.8
Hope this helps---ptaylor